OpenStudy (yananeedshelp):

what are the zeros f(x)=x^3-5x^2-6x?

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

\[ f(x)=x^3-5x^2-6x\] i think you can factor this one right?

OpenStudy (yananeedshelp):

Hi, can u help me please? :) and yes I need help

OpenStudy (misty1212):

each term has an \(x\) in it, "factor it out" and start with \[ x^3-5x^2-6x=0\\ x(x^2-5x-6)=0\]

OpenStudy (misty1212):

then you have to factor \(x^2-5x-6\) do you know how to do that?

OpenStudy (yananeedshelp):

no ma'am

OpenStudy (misty1212):

think of two numbers that multiplied together give \(-6\) and when you add then you get \(-5\) since the product is negative, one of the numbers has to be positive, the other negative

OpenStudy (misty1212):

did you get them yet?

OpenStudy (yananeedshelp):

one second

OpenStudy (yananeedshelp):

-2 * 3=-6 -6+1=-5

OpenStudy (misty1212):

they have to be the same for both

OpenStudy (misty1212):

like \(-2\times 3=-6\) and \(-2+3=1\) so that doesn't work try a different pair

OpenStudy (yananeedshelp):

-6*1=-6 -6+1=-5

OpenStudy (misty1212):

right

OpenStudy (misty1212):

\[x^2-5x-6=(x-6)(x+1)\]

OpenStudy (misty1212):

so now you have \[x(x-6)(x+1)=0\] and solve by setting each factor equal to zero

OpenStudy (misty1212):

\[x=0\\ x-6=0\iff x=6\\ x+1=0\iff x=-1\]

OpenStudy (yananeedshelp):

so the zeros are 6 and -1 ?